# Global minimality of the Hopf map in the Faddeev–Skyrme model with large coupling constant

### v1

@@ 4. Relaxed form of the perturbed Faddeev–Skyrme energy
Consider a closed $2$-form $\alpha$, thus $\alpha=d\beta$ for some $\beta\in W^{1,2}({\mathbb{S}}^{3};\bigwedge^{1}T^{*}{\mathbb{S}}^{3})$ with $d^{*}\beta=0$. The lowest eigenvalue of the Hodge Laplacian $\Delta:=dd^{*}+d^{\ast}d$ on closed $2$-forms is $4$, and the corresponding eigenspace is given by
$\ker d\cap\ker(\Delta-4)=E_{0}^{+}\oplus E_{0}^{-}\,,$
cf. [17, Section 3] and (2.9). Therefore, we can ($L^{2}$-orthogonally) decompose
$*\beta=\alpha_{0}^{+}+\alpha_{0}^{-}+d\varphi\,,\ \text{where }\ d*\alpha_{0}^{\pm}=\pm 2\alpha_{0}^{\pm}\,,\ \ d\varphi\perp_{L^{2}}E_{0}^{+}\oplus E_{0}^{-}\,,$ (5.2)
Using (5.2) and (5.3), we find
$\mathcal{I}_{2}(\alpha)=4\int_{\mathbb{S}^{3}}|\alpha_{0}^{+}|^{2}+4\int_{\mathbb{S}^{3}}|\alpha_{0}^{-}|^{2}+\int_{\mathbb{S}^{3}}|d\psi|^{2}\,,$
and
$Q(\alpha)=\frac{1}{16\pi^{2}}\int_{\mathbb{S}^{3}}\langle*\beta,\alpha\rangle\\ =\frac{1}{8\pi^{2}}\int_{\mathbb{S}^{3}}|\alpha_{0}^{+}|^{2}-\frac{1}{8\pi^{2}}\int_{\mathbb{S}^{3}}|\alpha_{0}^{-}|^{2}+\frac{1}{16\pi^{2}}\int_{\mathbb{S}^{3}}\psi\wedge d\psi\,.$
Combining the two expressions above, we deduce
$\begin{split}\mathcal{I}_{2}(\alpha)-32\pi^{2}Q(\alpha)&=8\int_{\mathbb{S}^{3}}|\alpha_{0}^{-}|^{2}+\int_{\mathbb{S}^{3}}|d\psi|^{2}-2\int_{\mathbb{S}^{3}}\psi\wedge d\psi\\ &\geq 8\int_{\mathbb{S}^{3}}|\alpha_{0}^{-}|^{2}+\int_{\mathbb{S}^{3}}|d\psi|^{2}-\frac{1}{4}\int_{\mathbb{S}^{3}}|d\psi|^{2}-4\int_{\mathbb{S}^{3}}|\psi|^{2}\,.\end{split}$ (5.4)
Note moreover that the next eigenvalue of $\Delta$ on closed $2$-forms is 9, hence (5.3) (from which we deduce that $d^{*}\psi=0$ and $\int_{{\mathbb{S}}^{3}}\psi=0$ component-wise), implies
$\int_{\mathbb{S}^{3}}|d\psi|^{2}\geq 9\int_{\mathbb{S}^{3}}|\psi|^{2}\,.$
This, together with (5.4) and again (5.3) yields
$\mathcal{I}_{2}(\alpha)-32\pi^{2}Q(\alpha)$ $\geq 8\int_{\mathbb{S}^{3}}|\alpha_{0}^{-}|^{2}+\bigg{(}1-\frac{1}{4}-\frac{4}{9}\bigg{)}\int_{\mathbb{S}^{3}}|d\psi|^{2}$
$\geq\frac{11}{36}\int_{{\mathbb{S}}^{3}}|\alpha-2\alpha_{0}^{+}|^{2}\geq\frac{1}{4}\min_{\widetilde{\alpha}^{+}_{0}\in E_{0}^{+}}\int_{\mathbb{S}^{3}}|\alpha-\widetilde{\alpha}^{+}_{0}|^{2}\,,$
proving (5.1). ∎
Corollary 5.2.
There exists $\rho_{0}\in(0,\sqrt{2})$ and $c>0$ such that the following holds: for any closed form $\alpha\in L^{2}\big{(}\mathbb{S}^{3};\bigwedge^{2}T^{*}\mathbb{S}^{3}\big{)}$ with $Q(\alpha)>0$, if $\alpha^{+}_{0}$ is the $L^{2}$-orthogonal projection of $\alpha$ onto $E_{0}^{+}$, we have
$\mathcal{E}_{\rho}(\alpha)-\mathcal{E}_{\rho}(\alpha^{+}_{0})\geq c\big{\|}Q(\alpha)^{-\frac{1}{2}}\alpha-Q(\alpha^{+}_{0})^{-\frac{1}{2}}\alpha^{+}_{0}\big{\|}_{L^{2}({\mathbb{S}}^{3})}^{2}\,,$ (5.5)
whenever $0<\rho\leq\rho_{0}$.
Proof.
Due to the homogeneity of the estimate we can assume that $Q(\alpha)=Q(\alpha^{+}_{0})=1$. It is also enough to find $\rho_{0}\in(0,\sqrt{2})$ for which (5.5) holds, as then by the fact that
$\mathcal{E}_{\rho}(\alpha)-\mathcal{E}_{\rho}(\alpha^{+}_{0})=\mathcal{E}_{\rho_{0}}(\alpha)-\mathcal{E}_{\rho_{0}}(\alpha^{+}_{0})+(\rho^{-2}-\rho_{0}^{-2})\left(\mathcal{I}_{2}(\alpha)-\mathcal{I}_{2}(\alpha^{+}_{0})\right)\geq\mathcal{E}_{\rho_{0}}(\alpha)-\mathcal{E}_{\rho_{0}}(\alpha^{+}_{0})\,,$
see Proposition 5.1, the same estimate holds for all $0<\rho\leq\rho_{0}$.
Let $\varepsilon_{0}:=\varepsilon_{0}(1)>0$ be the parameter provided by Proposition 4.2 with the choice $\widetilde{\rho}_{0}:=1$; we make this choice purely for convenience and there is nothing particular about it. Proposition 4.2 shows that if
$\big{\|}\alpha-\alpha^{+}_{0}\big{\|}_{L^{2}({\mathbb{S}}^{3})}\leq\varepsilon_{0}\,,$
and if we choose $\rho_{0}\in(0,1]$, then we have the desired inequality, thus we can suppose that $\|\alpha-\alpha^{+}_{0}\|_{L^{2}({\mathbb{S}}^{3})}\geq\varepsilon_{0}$. By the triangle inequality and the Cauchy-Schwarz inequality, we have
$\big{|}\mathcal{I}_{1}(\alpha)-\mathcal{I}_{1}(\alpha^{+}_{0})\big{|}\leq\int_{\mathbb{S}^{3}}|\alpha-\alpha^{+}_{0}|\leq\sqrt{2}\pi\|\alpha-\alpha^{+}_{0}\|_{L^{2}({\mathbb{S}}^{3})}\,,$
so that by Proposition 5.1 we find
$\mathcal{E}_{\rho_{0}}(\alpha)-\mathcal{E}_{\rho_{0}}(\alpha^{+}_{0})$ $\geq-2\sqrt{2}\pi\|\alpha-\alpha^{+}_{0}\|_{L^{2}({\mathbb{S}}^{3})}+\frac{\rho_{0}^{-2}}{4}\|\alpha-\alpha^{+}_{0}\|_{L^{2}({\mathbb{S}}^{3})}^{2}$
$\geq\bigg{(}\frac{\rho_{0}^{-2}}{4}-\frac{2\sqrt{2}\pi}{\varepsilon_{0}}\bigg{)}\|\alpha-\alpha^{+}_{0}\|_{L^{2}({\mathbb{S}}^{3})}^{2}\,.$
Hence the assertion of the corollary follows if we choose, e.g.$\rho_{0}:=\frac{1}{4}\big{(}\frac{\varepsilon_{0}}{\sqrt{2}\pi}\big{)}^{\frac{1}{2}}$. ∎
To obtain the uniqueness statement in Theorem 1.1 we will also need the following:
@@
$u=\psi\circ h\,.$ (5.9)
Indeed, since $h\colon{\mathbb{S}}^{3}\to{\mathbb{S}}^{2}$ is onto, for every $y\in\mathbb{S}^{2}$ we can find $x\in\mathbb{S}^{3}$ such that $h(x)=y$ and then we set $\psi(y):=u(x)$. This is well-defined, i.e., for $\mathcal{H}^{3}$-a.e. $x\in{\mathbb{S}}^{3}$, if $x^{\prime}\in{\mathbb{S}}^{3}$ is such that $h(x)=h(x^{\prime})$, so that $x,x^{\prime}$ belong to the same ${\mathbb{S}}^{1}$-fiber of $h$, then $u(x)=u(x^{\prime})$. For this, observe that by the general slicing properties of Sobolev functions, $u$ is absolutely continuous on a generic ${\mathbb{S}}^{1}$-fiber, and the claim follows from (5.6). That $\psi\in W^{1,2}({\mathbb{S}}^{2};{\mathbb{S}}^{2})$ follows trivially by the fact that $u\in W^{1,2}({\mathbb{S}}^{3};{\mathbb{S}}^{2})$ and that $h\colon{\mathbb{S}}^{3}\to{\mathbb{S}}^{2}$ is a $C^{\infty}$-submersion, cf. [21, Theorem 4.29] for the smooth case. In particular, we can pick smooth local coordinates $(\widetilde{x}_{1},\widetilde{x}_{2},\widetilde{x}_{3})$ on $\mathbb{S}^{3}$ such that $h(\widetilde{x})=(\widetilde{x}_{1},\widetilde{x}_{2})\in\mathbb{S}^{2}$ and we get the local expression $\psi(\widetilde{x}_{1},\widetilde{x}_{2})=u(\widetilde{x}_{1},\widetilde{x}_{2},\widetilde{x}_{3})$, hence $\psi$ has the same regularity as $u$.
@@
$\psi\in W^{1,2}({\mathbb{S}}^{2};{\mathbb{S}}^{2})\text{ is such that }d\psi^{t}d\psi=\mathrm{Id}_{T{\mathbb{S}}^{2}}\,\text{ with }\mathrm{deg}(\psi)=\pm 1\,.$ (5.10)
It follows that $\psi\equiv O\mathrm{id}_{{\mathbb{S}}^{2}}$ for some $O\in O(3)$, as can easily be inferred from Liouville’s theorem, see for instance [15, Theorem 1.1]. Finally, using the equivariance of the Hopf map exactly as at the end of the proof of [17, Proposition 4.2], there exists $R\in\textup{O}(4)$ such that
$u=\psi\circ h=h\circ R\,,$ (5.11)
@@
Proof of Theorem 1.1.
For $\rho_{0}>0$ as in Corollary 5.2 and any $\rho\in(0,\rho_{0}]$, Corollary 4.1, (2.16) and the scaling-invariant nature of (5.5) imply that
$\mathcal{FS}_{\rho}(u)-\mathcal{FS}_{\rho}(h)\geq\mathcal{E}_{\rho}(u^{*}\omega_{{\mathbb{S}}^{2}})-\mathcal{E}_{\rho}(h^{*}\omega_{{\mathbb{S}}^{2}})\geq c\min_{\widetilde{\alpha}^{+}_{0,1}\in E^{+}_{0,1}}\|u^{*}\omega_{{\mathbb{S}}^{2}}-\widetilde{\alpha}^{+}_{0,1}\|^{2}_{L^{2}({\mathbb{S}}^{2})}\geq 0\,,$ (5.12)
proving the minimality of $h$ among maps in $\mathcal{U}_{\mathcal{FS}}$. If $\widetilde{u}$ is another minimizer in this homotopy class, then equality holds in (5.12), i.e., $\widetilde{u}$ is horizontally weakly conformal by Corollary 4.1 and $\widetilde{u}^{*}\omega_{{\mathbb{S}}^{2}}\in E^{+}_{0,1}$, hence $|\widetilde{u}^{*}\omega_{{\mathbb{S}}^{2}}|=4$, cf. (4.11). By the invariance of the energy under rotations, using (2.12), (2.16) and (2.11), we can find $\widetilde{R}\in SO(4)$ so that the map $v:=\widetilde{u}\circ\widetilde{R}\in\mathcal{U}_{\mathcal{FS}}$, which is also a minimizer and horizontally weakly conformal, is such that $Q(v)=1$ and $v^{*}\omega_{{\mathbb{S}}^{2}}=h^{*}\omega_{{\mathbb{S}}^{2}}$. The conclusion then follows from Proposition 5.3. ∎
@@ Acknowledgments
The first two authors would also like to thank the Institute for Applied Mathematics of the University of Bonn, as well as the Hausdorff Center for Mathematics in Bonn, for their hospitality during the week 14.04.2025-18.04.2025, during which this work was initiated.

### v2

@@ 4. Relaxed form of the perturbed Faddeev–Skyrme energy
[…]
$\mathcal{FS}_{\rho}(u)-\mathcal{FS}_{\rho}(h)\geq\mathcal{E}_{\rho}(u^{*}\omega_{{\mathbb{S}}^{2}})-\mathcal{E}_{\rho}(h^{*}\omega_{{\mathbb{S}}^{2}})\geq 0\,.$ (4.32)
The characterization of the equality cases also follows from Proposition 4.2. ∎
5. Proof of uniqueness in Theorem 1.1
The uniqueness statement in Theorem 1.1 will follow from the following:
@@
$u=\psi\circ h\,.$
Indeed, since $h\colon{\mathbb{S}}^{3}\to{\mathbb{S}}^{2}$ is onto, for every $y\in\mathbb{S}^{2}$ we can find $x\in\mathbb{S}^{3}$ such that $h(x)=y$ and then we set $\psi(y):=u(x)$. This is well-defined, i.e., for $\mathcal{H}^{3}$-a.e. $x\in{\mathbb{S}}^{3}$, if $x^{\prime}\in{\mathbb{S}}^{3}$ is such that $h(x)=h(x^{\prime})$, so that $x,x^{\prime}$ belong to the same ${\mathbb{S}}^{1}$-fiber of $h$, then $u(x)=u(x^{\prime})$. For this, observe that by the general slicing properties of Sobolev functions, $u$ is absolutely continuous on a generic ${\mathbb{S}}^{1}$-fiber, and the claim follows from (5.1). That $\psi\in W^{1,2}({\mathbb{S}}^{2};{\mathbb{S}}^{2})$ follows trivially by the fact that $u\in W^{1,2}({\mathbb{S}}^{3};{\mathbb{S}}^{2})$ and that $h\colon{\mathbb{S}}^{3}\to{\mathbb{S}}^{2}$ is a $C^{\infty}$-submersion, cf. [lee2003smooth, Theorem 4.29] for the smooth case. In particular, we can pick smooth local coordinates $(\widetilde{x}_{1},\widetilde{x}_{2},\widetilde{x}_{3})$ on $\mathbb{S}^{3}$ such that $h(\widetilde{x})=(\widetilde{x}_{1},\widetilde{x}_{2})\in\mathbb{S}^{2}$ and we get the local expression $\psi(\widetilde{x}_{1},\widetilde{x}_{2})=u(\widetilde{x}_{1},\widetilde{x}_{2},\widetilde{x}_{3})$, hence $\psi$ has the same regularity as $u$.
@@
$\psi\in W^{1,2}({\mathbb{S}}^{2};{\mathbb{S}}^{2})\text{ is such that }d\psi^{t}d\psi=\mathrm{Id}_{T{\mathbb{S}}^{2}}\,\text{ with }\mathrm{deg}(\psi)=\pm 1\,.$
It follows that $\psi\equiv O\mathrm{id}_{{\mathbb{S}}^{2}}$ for some $O\in O(3)$, as can easily be inferred from Liouville’s theorem, see for instance [hirsch2022note, Theorem 1.1]. Finally, using the equivariance of the Hopf map exactly as at the end of the proof of [isobe2008, Proposition 4.2], there exists $R\in\textup{O}(4)$ such that
$u=\psi\circ h=h\circ R\,$,
@@
Completion of proof of Theorem 1.1.
If $\widetilde{u}$ is another minimizer in this homotopy class then equality holds in (4.32), i.e., $\widetilde{u}$ is horizontally weakly conformal by Corollary 4.1 and $\widetilde{u}^{*}\omega_{{\mathbb{S}}^{2}}\in E^{+}_{0,1}$, hence $|\widetilde{u}^{*}\omega_{{\mathbb{S}}^{2}}|=4$, cf. (4.9). By the invariance of the energy under rotations, using (2.12), (2.16) and (2.11), we can find $\widetilde{R}\in SO(4)$ so that the map $v:=\widetilde{u}\circ\widetilde{R}\in\mathcal{U}_{\mathcal{FS}}$, which is also a minimizer and horizontally weakly conformal, is such that $Q(v)=1$ and $v^{*}\omega_{{\mathbb{S}}^{2}}=h^{*}\omega_{{\mathbb{S}}^{2}}$. The conclusion then follows from Proposition 5.1. ∎
@@ Acknowledgments
The first two authors would also like to thank the Institute for Applied Mathematics of the University of Bonn, as well as the Hausdorff Center for Mathematics in Bonn, for their hospitality during the week 14.04.2025-18.04.2025, when this work was initiated. XL also thanks the Erwin Schrödinger International Institute for Mathematics and Physics (ESI), where part of this work was conducted during the 2025 thematic program on Free Boundary Problems.
Statements
Conflict of interest statement: The authors declare no conflict of interest.
Data availability statement: No data were used in the preparation of this manuscript.